The sum of the squares of three distinct real no.s which are in G.P. is S^2. If their sum is alphaS , show that alpha^2 belongs to open interval from 1/3 to 1 union open interval from 1 to 3.

Grade:11

1 Answers

Vikas TU
7 years ago
a + b +c = kS
here k is alpha.
and 
a^2 + b^2 + c^2 = S^2
and also,
b^2 = ac
Now U have three eqns. solve them and get a,b and c values and then think about the range of alpha simply.
but putting those values in any eqn to get the range of alpha that is k.